# Grid-based Multi-Robot Coverage Path Planning

### v1

@@ III-A Description of the Multi-Robot System
We consider a set $I=\{1,2,...,k\}$ of $k$ homogeneous and holonomic robots that operate in a four-way connected undirected 2d grid graph $G=(V,E)$, where $V$ and $E$ are the vertex set and edge set, respectively. Each vertex $v\in V$ is connected by edges to its top, bottom, left, and right neighbors if they exist. Each edge $e\in E$ is associated with a non-negative weight $w_{e}$, and we simply set $w_{e}=1$ for every $e\in E$ when considering an unweighted graph $G$.
We represent the robot state of arriving vertex $v\in V$ with a heading $\psi$ at time $t$ as $\mathbf{x}=(v,t,\psi)$. As graph $G$ is a four-way connected grid graph, we restrict the heading $\psi\in\{\text{E},\text{N},\text{W},\text{S}\}$ to East (E), North (N), West (W), and South (S). We assume the system has perfect knowledge about the robot states and put our focus of this paper on the time-optimal planning problems given in the next two subsections. Each robot $i$ is initialized in the state $(r_{i},0,\text{N})$, where $r_{i}\in V$ is called the root vertex given as an input. The robots can take the following three types of actions to transit from a state $\mathbf{x}=(v,t,\psi)$ to another state $\mathbf{x}^{\prime}=(v^{\prime},t^{\prime},\psi^{\prime})$:
  • Wait action: a robot waits at vertex $v$ for a time period $t^{\prime}-t>0$ where $v=v^{\prime}$ and $\psi=\psi^{\prime}$.
  • Turn action: a robot turns at vertex $v=v^{\prime}$ from $\psi$ to $\psi^{\prime}\neq\psi$ at the time cost of $t^{\prime}-t=c_{t}(\psi,\psi^{\prime})>0$, where $c_{t}$ is a function defined as $c_{t}(\psi,\psi^{\prime})=C\cdot\frac{|\psi^{\prime}-\psi|}{90^{\circ}}$. The constant $C$ is specified depending on the robots.
  • Move action: a robot moves from vertex $v$ to vertex $v^{\prime}\neq v$ via $e=(v,v^{\prime})\in E$ at the time cost of $t^{\prime}-t=w_{e}>0$.
In the rest of the paper, we will use tuple $(G,I,R)$ to denote the above Multi-Robot System (MRS) with a set $I$ of $k$ robots that live on the graph $G=(V,E)$ and are initialized in the set $R=\{r_{i}\}_{i=1}^{k}\subseteq V$ of root vertices. We aim to generate a set of plans for robots in the MRS to solve the problems in Sec. III-B, where a plan for any robot is defined as follows:
Definition 1 (Plan).
A plan $P=(\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{|P|})$ for the robot $i\in I$ is a finite ordered sequence of states such that $\mathbf{x}_{1}=(r_{i},0,\text{N})$, $\mathbf{x}_{|P|}=(r_{i},t_{\text{final}},\cdot)$, and every pair of adjacent states are in transit through wait, turn, or move actions.
The above definition follows the cover with return coverage problem setting [zheng2010multirobot], where the robots must start from and return to their corresponding initial vertices. The time cost $c(P)=t_{\text{final}}$ of plan $P$ is obtained from its last state $\mathbf{x}_{|P|}$. For convenience, we use $P[j]$ to represent the $j$-th state $\mathbf{x}_{i}$ of any plan $P$; when $j$ is negative, $P[j]$ represents the reversed $|j|$-th state $\mathbf{x}_{|P|+j+1}$ of $P$. We further denote the set of unique vertices visited by $P$ as $V(P)\subseteq V$, and the list of (potentially duplicated) edges used by $P$ as $E(P)$. To take the inter-robot conflict avoidance into consideration of planning, we formally define the conflict between any pair of plans as follows:
Definition 2 (Conflict).
Given any two plans $P_{a}$ and $P_{b}$, a conflict on vertex $v\in V$ occurs if there exists two states $P_{a}[i]=(v,t_{i},\cdot),P_{a}[i+1]=(\cdot,t^{\prime}_{i},\cdot)$ and two states $P_{b}[j]=(v,t_{j},\cdot),P_{b}[j+1]=(\cdot,t^{\prime}_{j},\cdot)$, such that the two time intervals $[t_{i},t^{\prime}_{i})$ and $[t_{j},t^{\prime}_{j})$ intersects with each other.
Our definition of conflict is more conservative than that in most MAPF algorithms. Specifically, we assume that a vertex is blocked not only when it is occupied by a robot, as is typical in the MAPF literature [sharon2015conflict], but also when a robot is moving out of it. By checking for time intervals during which other robots occupy a vertex, we can implicitly identify all conflict types outlined in the literature [stern2019multi], including vertex conflicts, edge conflicts, following conflicts, and swapping conflicts. These conflicts are typically defined for unweighted graphs, where time is discretized into uniform intervals for the MRS, allowing their actions to be aligned step by step. In addition, we adopt the conservative assumption of wellformness as defined in Definition 3 following the approach of [ma2019searching], which ensures that PBS-like algorithms are complete (i.e., always produce a valid solution) for various MAPF variants.
Definition 3 (Wellformness).
In a wellformed MRS $(G,I,R)$, each robot $i\in I$ can wait indefinitely at its root vertex $r_{i}\in V$ without blocking any other robot.
III-B Grid-based Multi-Robot Coverage Path Planning
For an MRS $(G,I,R)$ described above, we formally define the MCPP problem for the MRS as below:
@@ III-B Grid-based Multi-Robot Coverage Path Planning
Find a set $\mathcal{P}=\{P_{i}\}_{i=1}^{k}$ of plans for the set $I$ of robots such that $\bigcup_{i=1}^{k}V(P_{i})=V$ and there is no conflict between any pair of plans in $\mathcal{P}$.
Figure 2: The proposed algorithmic pipeline for MCPP (Problem 1) operates by first solving Relaxed-MCPP (Problem 2) to obtain a set of coverage paths, with which as input to further solve Deconflicted-MCPP (Problem 3) to obtain a set of plans, resulting the final set of action sequences for the robots.
We use makespan metric to evaluate the quality of an MCPP solution, denoted as $\tau(\mathcal{P})=\max\{c(P_{1}),c(P_{2}),...,c(P_{k})\}$. We emphasize that our formulation of the MCPP problem, as defined in Problem 1, differs from existing approaches in three key aspects: (1) it directly uses any edge-weighted graph as input, rather than the hypergraph of the graph to be covered; (2) it incorporates turning costs into the overall plan cost, not just traversal costs; and (3) it accounts for inter-robot conflicts. These distinctions significantly increase the difficulty of finding an optimal conflict-free solution, as we will demonstrate in Sec. III-C, where we show that optimally solving its two sub-problems are already NP-hard. To enhance the computational tractability of optimally solving Problem 1, we decompose it into two sub-problems. The first sub-problem addresses a relaxed version of MCPP by finding a set of paths that cover all vertices in $G$. These paths are specifically defined for the robots in the MRS $(G,I,R)$ as follows:
Definition 4 (Path).
A path $\pi=(v_{1},v_{2},...,v_{|\pi|})$ for the robot $i\in I$ is a finite ordered sequence of vertices such that $v_{1}=v_{|\pi|}=r_{i}$ and $(v_{j-1},v_{j})\in E$ for each $j=2,3,...,|\pi|$.
Similarly to the notation for a plan, we use $V(\pi)$, $E(\pi)$, and $\pi[j]$ to denote the set of unique vertices in $\pi$, the list of edges used by $\pi$, and the $j$-th vertex in $\pi$, respectively. The relaxed version of Problem 1 for MRS $(G,I,R)$, which corresponds to the classic MCPP problem without considering inter-robot conflicts as discussed in [tang2024large], is formally defined as follows:
Problem 2 (Relaxed-MCPP).
@@
After finding a set $\Pi$ of paths for Problem 2, we use it as input to solve the second sub-problem, which aims to construct a set $\mathcal{P}$ of conflict-free plans for Problem 1. Formally, the second sub-problem for MRS $(G,I,R)$ is defined as below:
Problem 3 (Deconflicted-MCPP).
Given a Problem 2 solution $\Pi=\{\pi_{i}\}_{i=1}^{k}$, find a set $\mathcal{P}=\{P_{i}\}_{i=1}^{k}$ of plans such that (1) there is no conflict between any pair of plans in $\mathcal{P}$; (2) for each $i\in I$, there exists a subsequence $(P_{i}[l_{1}],P_{i}[l_{2}],...,P_{i}[l_{|\pi_{i}|}])$ of $P_{i}$ where each $P_{i}[l_{j}]=(\pi_{i}[j],\cdot,\cdot)$ and each $l_{j}<l_{j+1}$.
In Problem 2, the paths from set $\Pi$ essentially disregard inter-robot conflicts and assume instantaneous heading changes between vertices, without considering intermediate turning actions for the robots. Consequently, for each robot $i\in I$, we can compute a cost $c(\pi_{i})$ as a lower bound of the true plan cost $c(P_{i})$ that accounts for conflicts as follows:
$c(\pi_{i})=\sum_{e\in E(\pi_{i})}w_{e}+\sum_{j=2}^{|\pi_{i}|}c_{t}(\psi_{j-1},\psi_{j})\leq c(P_{i})$ (1)
where each heading $\psi_{j}\in\{E,N,W,S\}$ is determined by the relative position between $\pi_{i}[j-1]$ and $\pi_{i}[j]$. As a lower bound of $c(P_{i})$, $c(\pi_{i})$ can guide the solution of Problem 2 by optimizing the makespan $\tau(\Pi)=\max\{c(\pi_{1}),c(\pi_{2}),...,c(\pi_{k})\}$. Ultimately, this approach aims to construct a makespan-minimized plan for Problem 1 by solving Problem 3. As shown in Fig. 2, in this paper, we propose to solve MCPP by sequentially solving Relaxed-MCPP and Deconflicted-MCPP.
III-C Theoretical Analysis
We show that optimally solving each sub-problem of MCPP is NP-hard, necessitating our time-efficient suboptimal algorithms in Sec. V and VI to solve them separately.
@@ III-C Theoretical Analysis
It is NP-hard to find an optimal Relaxed-MCPP solution with a minimized makespan.
Proof 1.
Inspired by [zheng2010multirobot], we prove Theorem 1 by reducing the strongly NP-hard 3-Partition problem (3-PP) [gary1979computers] to the decision version of Relaxed-MCPP, that is, determining whether Relaxed-MCPP has a solution with its makespan larger than a given positive number. Specifically, we consider a positive integer $B$ and a set $S=\{a_{i}\}_{i=1}^{3k}$ of positive integers with each $a_{i}\in(\frac{B}{4},\frac{B}{2})$ and $\sum_{i=1}^{3k}a_{i}=kB$, the objective of the 3-PP is to determine whether we can evenly partition $S$ into $k$ subsets $S_{1},S_{2},...,S_{k}$ of indices. We now construct a Relaxed-MCPP instance with the gird graph as shown in Fig. 3, where there is a horizontal “corridor” with $4k$ vertices $r_{n},r_{n-1},...,r_{1},v_{1},v_{2},...,v_{3k}$ at the center and $3k$ “tunnel” vertices $u_{1},u_{3},...,u_{3k}$ each connecting $v_{i}$. If index $i$ is odd, $u_{i}$ is on the top of $v_{i}$ otherwise on the bottom. The root vertex of robots $1,2,...,k$ are set to $r_{1},r_{2},...,r_{k}$, respectively. Each edge connecting corridor vertices have a weight of 1, whereas each tunnel vertex connecting the corridor vertex $v_{i}$ have a weight of $4ka_{i}$. As the instance construction can be completed in polynomial time, we now prove the claim that the makespan of the Relaxed-MCPP instance is at most $28kB+8k-$ 2 iff $S$ can be evenly partitioned into $k$ subsets:
$\bullet$ “If” direction: For each subset $S_{i}$ of the even partition, we know that $\sum_{i\in S_{j}}a_{i}=B$. Let the robot $j$ cover each tunnel vertex $u_{i}$ where $i\in S_{j}$, then its time cost is $\sum_{i\in S_{j}}4ka_{i}\times 2=8kB$ to cover the tunnel vertices and at most $(4k-1)\times 2$ to traverse through the corridor (from $r_{k}$ to $v_{3k}$). Combining both we have the time cost of each robot is at most $28kB+8k-$ 2 and thus the makespan is at most $28kB+8k-$ 2.
Figure 3: A Relaxed-MCPP instance reducible from the 3-Partition problem.
$\bullet$ “Only if” direction: We define $S_{j}$ to be the set of indices such that each $u_{i}$ for $i\in S_{j}$ is first covered by robot $j$, resulting in a partition of $k$ subsets $S_{1},S_{2},...,S_{k}$. As each $a_{i}\in(\frac{B}{4},\frac{B}{2})$ and $|S|=3k$, we know that each robot $j$ covers exactly three tunnel vertices and thus its time cost is at least $\sum_{i\in S_{j}}(4ka_{i}\times 2)+3\times 2=8k\sum_{i\in S_{j}}a_{i}+6$. Since the makespan is at most $28kB+8k-$ 2, we know the time cost of each robot $j$ is at most $28kB+8k-$ 2. As $a_{i}$, $B$ and $k$ are all positive integers, we have $\sum_{i\in S_{j}}a_{i}\leq B+1-\frac{1}{k}\leq B$ by chaining the above two inequalities. Since $\sum_{j=1}^{k}\sum_{i\in S_{j}}a_{i}=\sum_{i=1}^{3k}a_{i}=kB$, we have $\sum_{i\in S_{j}}a_{i}=B$ must hold to satisfy $\sum_{i\in S_{j}}a_{i}\leq B$, and thus $S_{1},S_{2},...,S_{k}$ is an even partition.
@@
It is NP-hard to find an optimal Deconflicted-MCPP solution with a minimized makespan.
Proof 2.
The OMG-MAPF problem (see Sec. II-B) was proved to be NP-hard with a makespan minimization objective [mouratidis2024fools], which can be viewed as a special case of Deconflicted-MCPP operating on unweighted graphs (i.e., $w_{e}=1$ for each $e\in E$). Therefore, it is at least NP-hard to find an optimal Deconflicted-MCPP solution with a minimized makespan.
IV Single-Robot Coverage Path Planning
In this section, we address the single-robot CPP problem, which is a specific case of MCPP (Problem LABEL:problem:_mcpp) when $k=|I|=1$ in instance $(G,I,R)$. Solving CPP efficiently is fundamental to our local search framework for MCPP in Sec. V. We propose Extended-STC (ESTC) to solve CPP, which builds on the STC algorithm for offline CPP and Full-STC [gabriely2002spiral] for online CPP. Both STC and Full-STC treat $G$ as a vertex-weighted graph to evaluate the traversal cost, where the weight of each vertex $v\in V$ depends on a hypervertex that can be decomposed into $v$ to reserve the solution optimality. However, this special treatment does not fully account for the actual traversal time cost incurred by a robot, which should be more naturally captured by the edges the robots traversed.
IV-A Extended Spanning Tree Coverage (ESTC)
ESTC requires constructing an auxiliary hypergraph $H=(V_{h},E_{h})$ from graph $G$ by contracting neighboring vertices within each grid cell into a hypervertex and connecting them with hyperedges between adjacent hypervertices. As shown in Fig. 4, the cells, represented by the light grey squares, are formed by contracting black vertices from $V$ into hypervertices. The graph $G$ is contracted left to right and bottom to top by default during the hypergraph construction. For clarity, we let $\delta_{v}\in V_{h}$ denote the hypervertex contracted from $v\in V$ and $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$ denote the hyperedge connecting $\delta_{u}$ and $\delta_{v}$. A hypervertex $\delta$ is termed incomplete if it is formed from fewer than four vertices; otherwise, it is complete. A special case arises when only two diagonal vertices exist within a grid cell —– each diagonal vertex forms its own nonadjacent hypervertex since they are not directly connected. To define the hyperedge weights for $H$, we categorize each hyperedge $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$ based on the number of crossing edges between the corresponding vertices within the cells of $\delta_{u},\delta_{v}$. As illustrated in Fig. 4, a crossing edge is an edge $(u,v)\in E$ where $u$ and $v$ are contracted into $\delta_{u}$ and $\delta_{v}$, respectively. The hyperedge weight $w_{\varepsilon}$ is then defined as follows:
$w_{\varepsilon}=\begin{cases}w_{(u_{1},v_{1})}+x_{(u_{2},v_{2})}-w_{(u_{1},u_{2})}-w_{(v_{1},v_{2})},&x_{\varepsilon}=2\\ 2\cdot w_{(u_{1},v_{1})},&x_{\varepsilon}=1\end{cases}$ (2)
where $x_{\varepsilon}$ is the number of crossing edges in $\varepsilon$. When $x_{\varepsilon}=0$, we know that $(\delta_{u},\delta_{v})\notin E_{h}$ and thus do not need to define its edge weight. Intuitively speaking, $w_{\varepsilon}$ implicitly encodes the cost difference between two circumnavigating coverage paths without explicitly generating them — one includes the crossing edges in $\varepsilon$ and the other does not. By Theorem 4, we demonstrate that Eqn. (2) ensures that ESTC (Alg. 1) always finds an optimal circumnavigating coverage path.
Figure 4: Illustration of crossing edges in hyperedge $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$.
We describe the pseudocode of ESTC in Alg. 1. After the hypergraph construction [Line 1], ESTC reduces the CPP problem to finding the Minimum Spanning Tree (MST) of $H$. Specifically, we use Kruskal’s algorithm to find the MST [Line 1] as the hyperedge weight could be negative. Then akin to STC, ESTC computes the coverage path by circumnavigating each MST edge via a Depth-First-Search [Line 1]. While existing STC-based CPP and MCPP methods [zheng2010multirobot, tang2023mixed] assume all hypervertices are complete, ESTC integrates routing rules from Full-STC to handle situations where hypervertices are incomplete, which redirects the path to the other side of a spanning edge when it is infeasible to move along one side due to the nonadjacency of vertices within the incomplete hypervertices. This makes ESTC a general algorithm to be applied on an arbitrary grid graph without assuming that all the hypervertices are complete, which is essential to the local search framework for MCPP in Sec. V.

### v2

@@ III MCPP and Its NP-Hardness
We consider a set $I=\{1,2,\ldots,k\}$ of $k$ robots operating on a four-neighbor connected undirected 2D grid graph $G=(V,E)$, where $V$ is the set of vertices and $E$ is the set of edges connecting each vertex to its top, bottom, left, and right neighbors if they exist. We assign each edge $e\in E$ with a non-negative weight $w_{e}$ and simply set $w_{e}=1$ for every $e\in E$ when considering an unweighted graph $G$. We specify a set $R=\{r_{i}\}_{i=1}^{k}\subseteq V$ of root vertices, where each robot $i$ starts at vertex $r_{i}$. We use the tuple $(G,I,R)$ to denote an MCPP instance.
A path $\pi_{i}$ for robot $i$ is defined as a finite ordered sequence of vertices $(v_{1},v_{2},\ldots,v_{|\pi_{i}|})$ such that $v_{1}=v_{|\pi_{i}|}=r_{i}$  [^2] The requirement that each robot $i$ must start and end at vertex $r_{i}$ is also known as the cover-and-return setting later in [13], which naturally aligns with STC (and also our ESTC) algorithm in Sec. IV since its coverage path forms a “loop”. A post-processing procedure can be found in [12] to convert a STC-like path for the cover-without-return setting. and $(v_{j-1},v_{j})\in E$ for each $j=2,3,\ldots,|\pi_{i}|$. The cost of any path $\pi$ is defined as $c(\pi)=\sum_{e\in\pi}w_{e}$. For clarity, we let $V(\pi)$ and $\mathcal{E}(\pi)$ denote the set of unique vertices in $\pi$ and the multiset of edges in $\pi$ that possibly contains repeated edges, respectively. The MCPP problem is defined as follows:
@@
The above problem formulation aligns with the cover with return coverage setting [13], where each robot must start from and return to its given root vertex. The quality of an MCPP solution is evaluated using the makespan metric, denoted by $\theta(\Pi)=\max\{c(\pi_{1}),c(\pi_{2}),\ldots,c(\pi_{k})\}$.
Remark 1.
Following conventional formulations, our MCPP problem definition intentionally excludes inter-robot deconfliction (see Definition 1). However, practical deployment demands collision-free paths. Previous coupled approaches for multi-goal sequencing MAPF [46, 51], a variant of our problem, typically scale only to a few dozen goals (vertices to be covered) and often have low success rates. Our two-phase approach addresses these limitations by first computing an MCPP solution without enforcing collision-free paths (Sec.V) and then deconflicting the solution via MAPF techniques (Sec.VI). This decoupled approach not only scales efficiently but is also more generally applicable: Any MCPP method can be used in the first phase, and our post-processing procedure can subsequently handle inter-robot conflicts.
MCPP can be viewed as an extension of min-max m-TSP on grid graphs where multiple depots exist and each city is allowed to be visited more than once. Given that the Hamiltonian cycle problem on grid graphs is NP-complete [52], both (single-robot) CPP and thus MCPP are NP-hard, which can be derived through a straightforward reduction from the Hamiltonian cycle problem.
@@
It is NP-hard to find a coverage path with the smallest cost on an unweighted grid graph.
Proof.
The Hamiltonian cycle problem, known to be NP-complete on a grid graph $G=(V,E)$ [52], can be directly reduced to the decision version of CPP that asks to determine whether there is a coverage path on the same grid graph $G$ with a cost not exceeding $|V|$. ∎
We now examine MCPP for a more restricted class of grid graphs. Although the Hamiltonian cycle problem can be solved in polynomial time [53] on solid grid graphs, which are grid graphs without holes, the scenario changes when considering MCPP. Specifically, inspired by the NP-hardness proof for a specialized MCPP variant defined on a vertex-weighted grid graph [13], we extend this complexity result to our more broadly defined MCPP on solid grid graphs through the following theorem.
@@
It is NP-hard to find an MCPP solution with the smallest makespan, even on a solid unweighted grid graph.
Proof.
We prove by showing the NP-completeness of a decision version of MCPP, that is, determining whether a given MCPP instance has a solution with a makespan not exceeding a given positive integer. One can evaluate whether the makespan of a given solution exceeds the given positive integer in polynomial time, implying that the MCPP problem is in NP. We now reduce the MCPP problem to the NP-complete 3-Partition problem [54] in polynomial time. Specifically, given a positive integer $B$ and a set $S=\{a_{i}\}_{i=1}^{3k}$ of positive integers, where each $a_{i}\in(\frac{B}{4},\frac{B}{2})$ and $\sum_{i=1}^{3k}a_{i}=kB$, the problem of 3-Partition is to determine if there exists an evenly partition $S$ into $k$ subsets $S_{1},S_{2},\ldots,S_{k}$, each summing to $B$.
Given a 3-Partition instance, we construct an MCPP instance on an unweighted grid graph as shown in Fig. (3). This graph has a central horizontal “corridor” with $4k$ vertices $r_{k},r_{k-1},\ldots,r_{1},v_{1},v_{2},\ldots,v_{3k}$ and $3k$ vertical “tunnel”, each connected to a vertex $v_{i}$. The $i$-th tunnel contains $4ka_{i}$ vertices and extends $v_{i}$ upwards if $i$ is odd and downwards if even. Robots $1,2,\ldots,k$ start at their root vertices $r_{1},r_{2},\ldots,r_{k}$, respectively.
We now show that the makespan of the above MCPP instance is at most $28kB+8k-$ 2 if and only if $S$ can be evenly partitioned into $k$ subsets:
[…]
Figure 3: An MCPP instance reducible from the 3-Partition problem.
IV Extended Spanning Tree Coverage (ESTC)
In this section, we present our new ESTC paradigm that advances standard STC-based paradigms by addressing coverage problems on any given connected grid graph, even when the quadrant coarsening $\mathcal{H}$ is incomplete. We first describe how ESTC solves (single-robot) CPP, a special case of MCPP (Problem 1) when $k=|I|=1$. Solving CPP efficiently is fundamental to our local search framework for MCPP detailed in Sec. V. We then analyze the properties and time complexity of ESTC and introduce two local optimizations to improve ESTC solutions. We finally discuss how ESTC can be integrated into existing MCPP methods to effectively handle incomplete $\mathcal{H}$.
ESTC extends both the STC paradigm for offline CPP and the Full-STC paradigm [55] for online CPP, which operate on a quadrant coarsened grid $\mathcal{H}$ derived from $G$. These paradigms typically evaluate the traversal costs by decomposing the given hypervertex weights of $\mathcal{H}$ into vertex weights for $G$ to maintain solution optimality. However, this approach does not adequately capture the actual traversal costs incurred by a robot, which are better represented by the edges traversed, as defined in our formulation.
IV-A Algorithm Description
ESTC operates on a hypergraph graph $H=(V_{h},E_{h})$, derived from the grid graph $G=(V,E)$. As shown in Fig. (5), similar to standard STC-based paradigms, ESTC attempts to contract neighboring vertices of $V$ corresponding to a $2\times 2$ block of grid cells into a hypervertex of $V_{h}$. However, unlike the quadrant coarsened grid $\mathcal{H}$ used by standard STC-based paradigms, the hypergraph $H$ effectively addresses the disconnections caused by partially obstructed $2\times 2$ blocks. A hyperedge connects two hypervertices only if there is at least one adjacent vertex pair from these hypervertices. A special case in hypervertex construction occurs when a $2\times 2$ block contains only two diagonal vertices. In this case, each diagonal vertex forms its own hypervertex, resulting in two nonadjacent hypervertices (see Fig. (5)). A hypervertex $\delta$ is incomplete if it is formed by fewer than four vertices; otherwise, it is complete. The graph $G$ is contracted left to right and bottom to top by default during the hypergraph construction. For clarity, we let $\delta_{v}\in V_{h}$ denote the hypervertex contracted from $v\in V$ and $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$ denote the hyperedge connecting $\delta_{u},\delta_{v}$.
Figure 4: Illustration of the ESTC algorithm. (a) The initial local paths (black lines). (b) The resulting ESTC coverage path by applying the rerouting rule on each hyperedge of a hypergraph spanning tree. (c)(d) For each hyperedge (gray solid line) $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$, the rerouting rule adds $E^{+}_{\varepsilon}$ (red solid lines) to and removes $E^{-}_{\varepsilon}$ (blue dashed lines) from the local paths of $\delta_{u}$ and $\delta_{v}$, depending on the number of the crossing edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ of $\varepsilon$.
Figure 5: The coarsened grid $\mathcal{H}$ in STC and the hypergraph $H$ in ESTC (both in gray markers and lines) for an input grid graph (black circles). The former yields incomplete coverage while the latter ensures complete coverage.
We describe the pseudocode for ESTC in Alg. 1. ESTC begins with the hypergraph construction [Line 1] and finds the Minimum Spanning Tree (MST) of $H$. Specifically, ESTC employs Kruskal’s algorithm [56] to find the MST [Line 1] since the hyperedge weights (will be defined later) can be negative. It then initializes a self-connecting local path $\pi_{\delta}$ for each $\delta_{v}\in V_{h}$ [Line 1], as shown in Fig. (4a). Note that $\pi_{\delta}$ is the optimal coverage path for the vertices with respect to hypervertex $\delta$. Then, for each hyperedge $\varepsilon=(\delta_{u},\delta_{v})$ in the MST $M$, a rerouting rule is applied to connect $\pi_{\delta_{u}}$ and $\pi_{\delta_{v}}$ to construct a coverage path for the vertices in $\delta_{u}$ and $\delta_{v}$ [Lines 1-1]. The resulting ESTC coverage path $\pi$ is returned after all the local paths are connected one to another [Line 1], as illustrated in Fig. (4b).
Rerouting Rule: The rerouting rule aims to connect pairwise self-connecting local paths of a hyperedge. We categorize each hyperedge $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$ based on the number of crossing edges between the corresponding vertices within the blocks of $\delta_{u},\delta_{v}$. A crossing edge is an edge $(u,v)\in E$ where $u$ and $v$ are contracted into adjacent hypervertices $\delta_{u}$ and $\delta_{v}$, respectively. Denote $x_{\varepsilon}$ as the number of crossing edges for a hyperedge $\varepsilon$, we define two multisets for any hyperedge $\varepsilon$, $E^{+}_{\varepsilon}=\{(u_{1},v_{1}),(u_{2},v_{2})\}$ and $E^{-}_{\varepsilon}=\{(u_{1},u_{2}),(v_{1},v_{2})\}$ if $x_{\varepsilon}=2$; otherwise when $x_{\varepsilon}=1$, $E^{+}_{\varepsilon}=\{(u_{1},v_{1}),(v_{1},u_{1})\}$ and $E^{-}_{\varepsilon}=\emptyset$. Note that $x_{\varepsilon}=0$ implies $(\delta_{u},\delta_{v})\notin E_{h}$. As illustrated in Fig. (4c) and (4d), $E^{+}_{\varepsilon}$ and $E^{-}_{\varepsilon}$ are represented in red solid lines and blue dashed lines, respectively. The rerouting rule for every $\varepsilon$ is then defined to add $E^{+}_{\varepsilon}$ to connect the two local paths while removing $E^{-}_{\varepsilon}$. It ensures that each vertex has an even degree with the least vertex duplication, making the connected path a valid path. The above rerouting rule can be interpreted as a combination of the standard STC circumnavigating pattern when $x_{\varepsilon}=2$ and the Full-STC rerouting rule when $x_{\varepsilon}=1$. It is always feasible as the construction of $H$ guarantees no vertex within any incomplete hypervertex is isolated, thereby ensuring the path remains feasible and continuous across any grid graphs.